# The commutativity of cofinal types

### v1

@@ Abstract.
We developed the theory of deterministic ideals and present a systematic study of the pseudo-intersection property with respect to an ideal introduced in [3]. We apply this theory to prove that for any two ultrafilters $U,V$ on $\omega$, $U\cdot V\equiv_{T}V\cdot U$. Additionally, we analyze the class of ultrafilters which are Tukey above $\omega^{\omega}$. Our theory applies the study of the Tukey types of a general sum of ultrafilters, which, as evidenced by the results of this paper, can be quite complex. Finally, we introduce the hierarchy of $\alpha$-almost rapid ultrafilters. We prove that it is consistent for them to form a strictly wider class than the rapid ultrafilters, and give an example of a non-rapid $p$-point which is Tukey above $\omega^{\omega}$. This answers several questions from [2, 3].
The research of the first author was supported by the National Science Foundation under Grant No. DMS-2246703
@@ 0. Introduction
The Tukey order stands out as one of the most studied orders of ultrafilters [28, 14, 25, 10, 32, 2]. Its origins lie in the examination of Moore-Smith convergence, and it holds particular significance in unraveling the cofinal structure of the partial order $(U,\supseteq)$ of an ultrafilter. Formally, given two posets, $(P,\leq_{P})$ and $(Q,\leq_{Q})$ we say that $(P,\leq_{P})\leq_{T}(Q,\leq_{Q})$ if there is map $f:Q\rightarrow P$, which is cofinal, namely, $f^{\prime\prime}\mathcal{B}$ is cofinal in $P$ whenever $\mathcal{B}\subseteq Q$ is cofinal. Schmidt [33] observed that this is equivalent to having a map $f:P\rightarrow Q$, which is unbounded, namely, $f^{\prime\prime}\mathcal{A}$ is unbounded in $Q$ whenever $\mathcal{A}\subseteq P$ is unbounded in $P$. We say that $P$ and $Q$ are Tukey equivalent, and write $P\equiv_{T}Q$, if $P\leq_{T}Q$ and $Q\leq_{T}P$; the equivalence class $[P]_{T}$ is called the Tukey type or cofinal type of $P$.
@@
The Tukey order is also related to the Katovich problem. A systematic study of the Tukey order on ultrafilter on $\omega$ traces back to Isbell [19], later to Milovich [28] and Dobrinen and Todorcevic [14]. Lately, Dobrinen and Benhamou [2] extended this study to ultrafilters on cardinals greater than $\omega$. They found a connection to a parallel recent line of research at the level of measurable cardinals, known as the Galvin property studied by Abraham, Benhamou, Garti, Goldberg, Gitik, Hayut, Magidor, Poveda, Shelah and others [1, 15, 16, 6, 4, 5, 8, 7, 17, 9]. The Galvin property in one of its forms is equivalent to being Tukey top as shown essentially by Isbell (in different terminology). Moreover, being Tukey-top in the restricted class of $\sigma$-complete ultrafilters takes the usual studied forms of the Galvin property.
In this paper, we address the problem of commutativity of the Tukey types of Fubini products of ultrafilters $U,V$ on $\omega$ (Definition 1.1), denoted by $U\cdot V$. This problem was suggested in [3], and was already partially addressed:
  • Dobrinen and Todorchevic [14] proved that if $U,V$ are rapid $p$-points then $U\cdot V\equiv_{T}V\cdot U$.
@@
The main result of this paper is:
@@
For any ultrafilters $U_{1},U_{2},...,U_{n}$ on $\omega$, and any permutation $\sigma:\{1,...,n\}\rightarrow\{1,...,n\}$, $U_{1}\cdot U_{2}\cdot...\cdot U_{n}\equiv_{T}U_{\sigma(1)}\cdot...\cdot U_{\sigma(n)}$.
To do this we will, in fact, prove that
(*$(*)\ \ \ U_{1}\cdot U_{2}\cdot...\cdot U_{n}\equiv_{T}U_{1}\times U_{2}\times...\times U_{n}\times(U_{1}\cap...\cap U_{n})^{\omega}$
from which the commutativity follows straightforwardly.
Our main result stands in sharp contrast to the Rudin-Keisler ordering which is known not to be commutative with respect to Fubini product [^2] For example if $U,W$ are non-isomorphic Ramsey ultrafilters then $U\cdot W\not\equiv_{RK}W\cdot U$. Just otherwise, by a theorem of Rudin (see for example [20, Thm. 5.5], $U,W$ should be Rudin-Frolík (and therefore Rudin-Keisler) comparable, contradicting the RK-minimality of Ramsey ultrafilters.. On measurable cardinals, the situation is even more dramatic, due to a theorem of Solovey (see [20, Thm. 5.7]) if $U,W$ are $\kappa$-complete ultrafilters on $\kappa$ the $U\cdot W\equiv_{T}W\cdot U$ if and only if $W\equiv_{RK}U^{n}$ for some $n$ or vise versa. Recently, Goldberg [18] examined situations of commutativity with respect to several product operations of countable complete ultrafilters.
As a corollary of our main theorem, we conclude the following Tukey-structural results:
@@
The class of ultrafilters $U$ such that $U\cdot U\equiv_{T}U$ is upward closed with respect to the Tukey order.
The main idea that is used in the proof of the main result, as suggested by the equivalence (*)(*)( * ), is to analyze the cofinal types of ideals and filters connected to a given ultrafilter $U$. More specifically, we will exploit the idea of the pseudo-intersection property with respect to $I$ (Definition 2.2) which was introduced in [3] and was used to prove that Milliken-Taylor ultrafilters and generic ultrafilters [^3] For the definition of $I^{\otimes\alpha}$, see the paragraph before Fact 2.11. for $P(\omega^{\alpha})/\mathrm{fin}^{\otimes\alpha}$ satisfy $U\cdot U\equiv_{T}U$. In  §2, we provide a comprehensive study of this property. The main result of this section is
@@
Suppose that $\mathcal{A}$ is a discrete set of ultrafilters. Then for each $U\in\mathcal{A}$, $U$ has the $(\cap\mathcal{A})^{*}$-p.i.p.
@@
For complete directed ordered set $\mathbb{P}$, $\mathbb{P}$ is uniformly bellow [^5] See definition $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\kappa{\rangle})$ if and only if $\sum_{U}V_{\alpha}\geq_{T}\mathbb{P}$.
By putting more assumptions on the sequence of ultrafilters, we are able to get nicer results. One of them is that under the assumption that
@@
in which case we have that $\sum_{U}V_{n}\equiv_{T}U\times\prod_{n<\omega}V_{n}$. From this special case, we can recover the formula $U\cdot V\equiv_{T}U\times V^{\omega}$. Another set of assumptions we consider is when for each $n<\omega$, $V_{n}\cdot V_{n}\equiv_{T}V_{n}\geq_{T}V_{n+1}$. These assumptions imply that $\sum_{U}V_{n}$ is a strict greatest least lower bound of $\mathcal{B}(U,{\langle}V_{n}\mid n<\omega{\rangle})$. This shows that the cofinal type of $\sum_{U}V_{n}$ is much more complicated than the cofinal type of $U\cdot V$. We also provide an example (Prop. 1.21) where $U<_{T}\sum_{U}V_{n}<U\times\prod_{n\in X}V_{n}$ for every $X\in U$.
@@
this partial order appeared in the work of Molvice [28, Question 4.7], and Dobrinen and Todorcevic who proved the following [14, Thm. 35]:
@@
They then proved that rapid ultrafilters are above $\omega^{\omega}$ and deduced that rapid $p$-points satisfy $U\cdot U\equiv_{T}U$. This was lately improved by Benhamou and Dobrinenn [3, Thm 1.18] to the general setup of $I$-p.i.p.
@@
Taking $I=\mathrm{fin}$ reproduces the difficult part from Dobrinen and Todorcevic’s result [^6] Indeed, as pointed out by Dobrinen and Todorcevic, it is easy to see that $U\cdot U\geq_{T}\omega^{\omega}$, so $(1)\Rightarrow(2)$ is straightforward.. In  §4, we first note that in some sense the assumption that $U$ is $I$-p.i.p in the theorem above is not optimal.
@@
Let $U$ be any ultrafilter over a set $X$. If $U$ does not have the $I$-p.i.p then $U\geq_{T}\prod_{x\in X}I$.
In particular, if $U$ is not a $p$-point then automatically $U\geq_{T}\omega^{\omega}$. Hence, to understand the class of ultrafilters Tukey above $\omega^{\omega}$, it suffices to restrict our attention to $p$-points. Inside the class of $p$-point, it is known to be consistent that there are $p$-point which are not above $\omega^{\omega}$ (see [14]), and the class of rapid p-point is currently the largest subclass of $p$-points known to be above $\omega^{\omega}$. In the second part of Section 4, we enlarge this class by introducing the notion of $\alpha$-almost-rapid (Definition 4.11), which is a weakening of rapidness.
@@
  • In §3, we prove our main results.
@@
$[X]^{<\lambda}$ denotes the set of all subsets of $X$ of cardinality less than $\lambda$. Let $\mathrm{fin}=[\omega]^{<\omega}$, and $\mathrm{FIN}=\mathrm{fin}\setminus\{\emptyset\}$. For a collection of sets $(P_{i})_{i\in I}$ we let $\prod_{i\in I}P_{i}=\{f:I\rightarrow\bigcup_{i\in I}P_{i}\mid\forall i,\,f(i)\in P_{i}\}$. If $P_{i}=P$ for every $i$, then $P^{I}=\prod_{i\in I}P$. Given a set $X\subseteq\omega$, such that $|X|=\alpha\leq\omega$, we denote by ${\langle}X(\beta)\mid\beta<\alpha{\rangle}$ be the increasing enumeration of $X$. Given a function $f:A\rightarrow B$, for $X\subseteq A$ we let $f^{\prime\prime}X=\{f(x)\mid x\in X\}$, for $Y\subseteq B$ we let $f^{-1}Y=\{x\in X\mid f(x)\in Y\}$, and let $\operatorname{rng}(f)=f^{\prime\prime}A$. Given sets $\{A_{i}\mid i\in I\}$ we denote by $\biguplus_{i\in I}A_{i}$ the union of the $A_{i}$’s when the sets $A_{i}$ are pairwise disjoint. Two partially ordered set $\mathbb{P},\mathbb{Q}$ are isomorphic, denoted by $\mathbb{P}\simeq\mathbb{Q}$, if there is a bijection $f:\mathbb{P}\rightarrow\mathbb{Q}$ which is order-preserving.
@@ 1.1. Some basic definitions and facts
[…]
@@
Next, let us record some basic terminology and facts regarding cofinal types. Given two directed partially ordered sets $\mathbb{P},\mathbb{Q}$, the Cartesian product $\mathbb{P}\times\mathbb{Q}$ ordered pointwise, is the least upper bound of $\mathbb{P},\mathbb{Q}$ in the Tukey order (see [13]). It follows that $F\times G\leq_{T}F\cdot G$. More generally, for partially ordered sets $(\mathbb{P}_{i},\leq_{i})$ for $i\in I$, we denote by $\prod_{i\in I}(\mathbb{P}_{i},\leq_{i})$ to be the order over the underlining set $\prod_{i\in I}\mathbb{P}_{i}$ with the everywhere domination order, namely $f\leq g$ iff for all $i\in I$, $f(i)\leq_{i}g(i)$. If the order is clear from the context we omit it and just write $\prod_{i\in I}\mathbb{P}_{i}$. This is the case when $\mathbb{P}_{i}=U_{i}$ is a filter ordered by reversed inclusion of an ideal ordered by inclusion. If for every $i\in I$, $\mathbb{P}_{i}=\mathbb{P}$ we simply write $\mathbb{P}^{I}$.

### v2

# Commutativity of cofinal types

@@ Abstract.
We continue the study of the pseudo-intersection property with respect to an ideal introduced in [3]. Our theory applies to the study of the Tukey types of general sums of ultrafilters, which, as evidenced by the results of this paper, can be quite complex. It also applies to construct a large class of ultrafilter $\mathcal{C}$ over $\omega$ such that any two ultrafilters $U,V\in\mathcal{C}$ commute; that is, $U\cdot V\equiv_{T}V\cdot U$. The class $\mathcal{C}$ class contains most known cofinal types of ultrafilters on $\omega$. This is in sharp contrast to the Rudin-Keisler ordering. In the third part of this paper, we apply our results to study the class of ultrafilters Tukey above $\omega^{\omega}$. Specifically, we prove that ultrafilters without the $I$-p.i.p are always above $I^{\omega}$ and in particular non-$p$-points are Tukey above $\omega^{\omega}$. Finally, we introduce the hierarchy of $\alpha$-almost rapid ultrafilters. We prove that it is consistent for them to form a strictly wider class than the rapid ultrafilters, and give an example of a non-rapid $p$-point ultrafilter which is Tukey above $\omega^{\omega}$. This addresses and answers several questions from [2, 3, 14, 28].
Key words and phrases:
ultrafilter, $p$-point, rapid ultrafilter, Tukey order, cofinal type, Fubini product
2020 Mathematics Subject Classification:
03E02, 03E04, 03E05, 03E55, 06A06, 06A07
This research was supported by the National Science Foundation under Grant No. DMS-2346680
@@ 0. Introduction
The Tukey order stands out as one of the most studied orders of ultrafilters [28, 14, 25, 10, 32, 2]. Its origins lie in the examination of Moore-Smith convergence, and it holds particular significance in unraveling the cofinal structure of the partial order $(U,\supseteq)$ of an ultrafilter. Formally, given two posets, $(P,\leq_{P})$ and $(Q,\leq_{Q})$ we say that $(P,\leq_{P})\leq_{T}(Q,\leq_{Q})$ if there is map $f:Q\rightarrow P$, which is cofinal, namely, $f^{\prime\prime}\mathcal{B}$ is cofinal in $P$ whenever $\mathcal{B}\subseteq Q$ is cofinal. Schmidt [33] observed that this is equivalent to having a map $f:P\rightarrow Q$, which is unbounded, namely, $f^{\prime\prime}\mathcal{A}$ is unbounded in $Q$ whenever $\mathcal{A}\subseteq P$ is unbounded in $P$. We say that $P$ and $Q$ are Tukey equivalent, and write $P\equiv_{T}Q$, if $P\leq_{T}Q$ and $Q\leq_{T}P$; the equivalence class $P]_{T}$ is called the Tukey type or cofinal type of $P$.
@@
The Tukey order is also related to the Katovich problem. A systematic study of the Tukey order on ultrafilter over $\omega$, traces back to Isbell [19], later to Milovich [28] and Dobrinen and Todorcevic [14]. Lately, Benhamou and Dobrinen [2] extended this study to ultrafilters on cardinals greater than $\omega$. Over measurable cardinals, the Tukey order is connected to recent developments revolving the so-called Galvin property, studied by Abraham, Benhamou, Garti, Goldberg, Gitik, Hayut, Magidor, Poveda, Shelah and others [1, 15, 16, 6, 4, 5, 8, 7, 17, 9]; the Galvin property in one of its forms is equivalent to being Tukey-top as shown essentially by Isbell (in different terminology). Moreover, being Tukey-top in the restricted class of $\kappa$-complete ultrafilters takes the usual studied forms of the Galvin property.
In this paper, we address the problem of commutativity of the Tukey types of Fubini products of ultrafilters $U,V$ over $\omega$ (Definition 1.1), denoted by $U\cdot V$. This problem was suggested in [3], and was already partially addressed:
  • Dobrinen and Todorcevic [14] proved that if $U,V$ are rapid $p$-points then $U\cdot V\equiv_{T}V\cdot U$.
@@
Our goal is to study larger classes where commutativity holds, these are called commutative classes of ultrafilters. In that avenue, we prove the following:
@@
There is a commutative class of ultrafilter $\mathcal{C}$ which is closed under Fubini sums and includes all Tukey-top ultrafilters, $p$-points, stable ordered union ultrafilters, generic ultrafilters for $P(\omega)/I$ for a wide range of $I$’s.  [^2] II should be simple, and $I^{\omega}\equiv_{V}I$, so for example $I$ can be $\mathrm{fin}^{\otimes\alpha}$.
The commutative class described in the previous theorem includes most known ultrafilters on $\omega$, and in fact it is open whether this class consistently excludes an ultrafilter on $\omega$. The commutativity of cofinal types stands in sharp contrast to the Rudin-Keisler ordering which is known to be highly non commutative with respect to Fubini product [^3] For example if $U,W$ are non-isomorphic Ramsey ultrafilters then $U\cdot W\not\equiv_{RK}W\cdot U$. Just otherwise, by a theorem of Rudin (see for example [20, Thm. 5.5], $U,W$ should be Rudin-Frolík (and therefore Rudin-Keisler) comparable, contradicting the RK-minimality of Ramsey ultrafilters.. On measurable cardinals, the situation is even more dramatic, due to a theorem of Solovey (see [20, Thm. 5.7]) if $U,W$ are $\kappa$-complete ultrafilters on $\kappa$ the $U\cdot W\equiv_{T}W\cdot U$ if and only if $W\equiv_{RK}U^{n}$ for some $n$ or vise versa. Recently, Goldberg [18] examined situations of commutativity with respect to several product operations on countably complete ultrafilters.
The main idea that is used in the proof, is to analyze the cofinal types of ideals and filters connected to a given ultrafilter $U$. More specifically, we will exploit the idea of the pseudo-intersection property with respect to $I$ (Definition 2.2) which was introduced in [3] and was used to prove that Milliken-Taylor ultrafilters and generic ultrafilters [^4] For the definition of $I^{\otimes\alpha}$, see the paragraph before Fact 2.11. for $P(\omega^{\alpha})/\mathrm{fin}^{\otimes\alpha}$ satisfy $U\cdot U\equiv_{T}U$. In fact we use the theory developed here to prove some generalization of theorem regarding generic ultrafilters over $P(\omega^{\alpha})/\mathrm{fin}^{\otimes\alpha}$:
@@
Let $I$ be a deterministic $\sigma$-ideal [^5] A $\sigma$-ideal is an ideal such that $P(\omega)/I$ is $\sigma$-complete. Then for every $\alpha<\omega_{1}$, and any $V$-generic ultrafilter for $P(\omega^{\alpha})/I^{\otimes\alpha}$, $G\cdot G\equiv_{T}G$.
In  §2, we provide a comprehensive study of this property. The main result of this section is
@@
Suppose that $\mathcal{A}$ is a discrete set of ultrafilters. Then for each $\in\mathcal{A}, UU$ has the $(\bigcap\mathcal{A})^{*}$-p.i.p.
@@
For complete directed ordered set $\mathbb{P}$, $\mathbb{P}$ is uniformly below  [^7] See Definition 1.11. $\mathcal{B}(U,{\langle}V_{\alpha}\mid\alpha<\kappa{\rangle})$ if and only if $\sum_{U}V_{\alpha}\geq_{T}\mathbb{P}$.
By putting more assumptions on the sequence of ultrafilters, we are able to get nicer results. One of them is that under the assumption
@@
in which case we have that $\sum_{U}V_{n}\equiv_{T}U\times\prod_{n<\omega}V_{n}$. From this special case, we recover D. Milovich’s formula $U\cdot V\equiv_{T}U\times V^{\omega}$ (See Theorem 1.7). Another set of assumptions we consider is when for each $n<\omega$, $V_{n}\cdot V_{n}\equiv_{T}V_{n}\geq_{T}V_{n+1}$. These assumptions imply that $\sum_{U}V_{n}$ is a strict greatest least lower bound of $\mathcal{B}(U,{\langle}V_{n}\mid n<\omega{\rangle})$. This shows that the cofinal type of $\sum_{U}V_{n}$ is much more complicated than the cofinal type of $U\cdot V$. We also provide an example (Prop. 1.21) where $U<_{T}\sum_{U}V_{n}<U\times\prod_{n\in X}V_{n}$ for every $X\in U$.
@@
On the other hand, among analytic ideals, it is known to be minimal [37](See also [25, Thm 6.6]):
Theorem 0.2 (Todorcevic).
Suppose that $I$ is an analytic $p$-ideal, then either $I$ is countable generated or $I\geq_{T}\omega^{\omega}$.
This was later improved by Solecki and Todorcevic [36, Proposition 4.3] to show that if $I$ is analytic, not locally compact ideal, then $I\geq_{T}\omega^{\omega}$. This partial order came up in the work of Milovich who asked [28, Question 4.7] if there is an ultrafilter $U$ over $\omega$ such that $(U,\supseteq)\equiv_{T}\omega^{\omega}$. We will observe that this was basically answered in [36, Cor. 54]:
Theorem 0.3 (Solecki-Todorcevic).
Suppose that $D$ is an ordered separable metric space such that the predecessors of each element form a compact set, and $E$ is a basic  [^8] For the definition of basic see [36, §3]. analytic order such that $D\leq_{T}E$, then $D$ is analytic.
Later, Dobrinen and Todorcevic contributed a great deal to the understanding of this class, in particular, they proved the following [14, Thm. 35]:
@@
They also proved that rapid ultrafilters are above $\omega^{\omega}$ and deduced that rapid $p$-points satisfy $U\cdot U\equiv_{T}U$. This was lately improved by Benhamou and Dobrinenn [3, Thm 1.18] to the general setup of the $I$-p.i.p.
@@
Taking $I=\mathrm{fin}$ reproduces the difficult part from Dobrinen and Todorcevic’s result [^9] Indeed, as pointed out by Dobrinen and Todorcevic, it is easy to see that $U\cdot U\geq_{T}\omega^{\omega}$, so $(1)\Rightarrow(2)$ is straightforward.. In  §4, we first note that in some sense the assumption that $U$ has the $I$-p.i.p in the theorem above is not optimal.
@@
Let $U$ be any ultrafilter over a set $X$. If $U$ does not have the $I$-p.i.p then $U\geq_{T}I^{X}$.
In particular, we get the following corollary:
Corollary.
If $U$ is not a $p$-point ultrafilter over $\omega$ then $U\geq_{T}\omega^{\omega}$.
This last corollary is of the same flavor as Theorems 0.1,0.2. Hence if we drop the $p$-point assumption, then we also get $U\geq_{T}\omega^{\omega}$. Hence, if we are looking for examples for ideals that are not Tukey above $\omega^{\omega}$ and are not countably generated, we might as well restrict our attention to non-analytic $p$-ideals. This is closely related to [14, Question 42].
Inside the class of $p$-point, it is known to be consistent that there are $p$-points which are not above $\omega^{\omega}$ (see [14]), and the class of rapid p-point is currently the largest subclass of $p$-points known to be above $\omega^{\omega}$. In the second part of Section 4, we enlarge this class by introducing the notion of $\alpha$-almost-rapid (Definition 4.12), which is a weakening of rapidness.
@@
  • In §3, we prove our results regarding the commutative classes.
@@
$X]^{<\lambda}$ denotes the set of all subsets of $X$ of cardinality less than $\lambda$. Let $\mathrm{fin}=[\omega]^{<\omega}$, and $\mathrm{FIN}=\mathrm{fin}\setminus\{\emptyset\}$. For a collection of sets $(P_{i})_{i\in I}$ we let $\prod_{i\in I}P_{i}=\{f:I\rightarrow\bigcup_{i\in I}P_{i}\mid\forall i,\,f(i)\in P_{i}\}$. If $P_{i}=P$ for every $i$, then $P^{I}=\prod_{i\in I}P$. Given a set $X\subseteq\omega$, such that $|X|=\alpha\leq\omega$, we denote by ${\langle}X(\beta)\mid\beta<\alpha{\rangle}$ be the increasing enumeration of $X$. Given a function $f:A\rightarrow B$, for $X\subseteq A$ we let $f^{\prime\prime}X=\{f(x)\mid x\in X\}$, for $Y\subseteq B$ we let $f^{-1}Y=\{x\in X\mid f(x)\in Y\}$, and let $\operatorname{rng}(f)=f^{\prime\prime}A$. Given sets $\{A_{i}\mid i\in I\}$ we denote by $\biguplus_{i\in I}A_{i}$ the union of the $A_{i}$’s when the sets $A_{i}$ are pairwise disjoint. Two partially ordered set $\mathbb{P},\mathbb{Q}$ are isomorphic, denoted by $\mathbb{P}\simeq\mathbb{Q}$, if there is a bijection $f:\mathbb{P}\rightarrow\mathbb{Q}$ which is order-preserving.
@@ 1.1. Some basic definitions and facts
where $(A)_{x}=\{y\in Y_{x}\mid{\langle}x,y{\rangle}\in A\}$ is the $x^{\text{th}}$-fiber of $A$. If for every $x$, $G_{x}=G$ for some fixed $V$ over a set $Y$, then $F\cdot G$ is defined as $\sum_{F}G$, which is a filter over $X\times Y$. $F^{2}$ denotes the filter $F\cdot F$ over $X\times X$.
@@
Next, let us record some basic terminology and facts regarding cofinal types. Given two directed partially ordered sets $\mathbb{P},\mathbb{Q}$, the Cartesian product $\mathbb{P}\times\mathbb{Q}$ ordered pointwise, is the least upper bound of $\mathbb{P},\mathbb{Q}$ in the Tukey order (see [13]). It follows that $F\times G\leq_{T}F\cdot G$. More generally, for partially ordered sets $(\mathbb{P}_{i},\leq_{i})$ for $i\in I$, we denote by $\prod_{i\in I}(\mathbb{P}_{i},\leq_{i})$ to be the order over the underlining set $\prod_{i\in I}\mathbb{P}_{i}$ with the everywhere domination order, namely $f\leq g$ iff for all $i\in I$, $f(i)\leq_{i}g(i)$. If the order is clear from the context we omit it and just write $\prod_{i\in I}\mathbb{P}_{i}$. This is the case when $\mathbb{P}_{i}=U_{i}$ is a filter ordered by reversed inclusion of an ideal ordered by inclusion. If for every $\in I, i=\mathbb{P}_{i}$$\mathbb{P}$ we simply write $\mathbb{P}^{I}$.